A total investment of $9,700is made into two savings accounts. One account yields 9% simple interest and the other 10% simple interest. He earns a total of $935.00interest for the year. How much was invested in the 9% account?
Answer by Boreal(15235)

x=amount in 9%
9700-x=amount at 10%
interest= .09x+(9700-x)(.10)=935
0.09 + 970-(.10x)=935 by distributing
-0.01x=-35. multiply by $100
x=3500 @9%=$315
$6200 @10%=$620
They sum to $935